OpenStudy (annah.bee):

Help me please with this question? Is the function linear or nonlinear? y= 1 ---- x A. linear B. nonlinear

Vocaloid (vocaloid):

start by plugging in values when x = 1, what does y = ?

OpenStudy (annah.bee):

1/2 ?

OpenStudy (annah.bee):

@Vocaloid

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

y = 1/x = ?

OpenStudy (annah.bee):

1

Vocaloid (vocaloid):

good now what does y equal when x = 2?

OpenStudy (annah.bee):

1/2

OpenStudy (annah.bee):

but how could 1/2 be put on a graph?

Vocaloid (vocaloid):

you put it between 1 and 2...

Vocaloid (vocaloid):



Vocaloid (vocaloid):

ready to keep going?

OpenStudy (annah.bee):

yes

OpenStudy (annah.bee):

btw thank you so much for helping me

Vocaloid (vocaloid):

good, now what does y equal when x = 3?

OpenStudy (annah.bee):

1/3

Vocaloid (vocaloid):

good, so now we have x = 1, y = 1 x = 2, y = 1/2 x = 3, y = 1/3 is y changing at a constant rate?

OpenStudy (annah.bee):

i dont think so

Vocaloid (vocaloid):

good, so it's nonlinear

OpenStudy (annah.bee):

okay i see that now thank you so much your a life saver

Vocaloid (vocaloid):

this is what the graph looks like if you're curious definitely not a straight line